# SPECIAL RELATIVITY

### v11

@@
New material has been added in the Chapter 4 and Chapter 16. If you have any comments, please send an email to evgueni.saldin@desy.de.
@@ 1 Introduction
Only by solving the equations of motion in their fully covariant form can one correctly capture the coupling between Maxwell’s equations and particle trajectories in the lab frame. Solving problems related to covariant particle tracking presents two key challenges. First, conventional particle tracking within a single inertial frame is typically based on classical Newtonian mechanics, which does not incorporate Lorentz transformations. Second, there is considerable confusion regarding the combination of non-parallel velocities in special relativity. The standard textbook treatment of relativistic velocity transformation implicitly assumes that the $(x^{\prime},y^{\prime},z^{\prime})$ axes of a moving observer remain parallel to the $(x,y,z)$ axes of the lab frame. In other words, it presumes that both observers share a common three-dimensional space—a misconception. Textbooks often overlook the fact that in Lorentz coordinatization, two observers following different trajectories experience different three-spaces. Special relativity lacks an absolute notion of simultaneity, leading to a fundamental mixing of spatial and temporal coordinates. For one observer, spatial measurements inherently include a contribution from the time coordinate as perceived by another observer. Consequently, there is no well-defined concept of an instantaneous three-space.
@@
In Chapter 16, we explore which aspects of dynamical theory depend on the choice of synchronization convention and which remain independent. Our analysis highlights the distinction between the concepts of path and trajectory. A path is a purely geometric notion, representing the spatial course of a particle without reference to time. In contrast, a trajectory provides more information, as it specifies not only the particle’s spatial position but also the corresponding time instant. The path has a precise, objective meaning. For example, in a magnetic field, the curvature radius of the path—and consequently, the three-momentum—is convention-invariant and thus objectively defined. However, the trajectory $\vec{x}(t)$ is convention-dependent, reflecting the inherent conventionality of velocity, and therefore lacks an exact objective meaning. While dynamical theory involves the concept of a particle’s trajectory, direct verification of it is unnecessary. Instead, the trajectory serves as a theoretical tool in the analysis of electrodynamics problems.
@@
In Chapter 18, we present a critical reexamination of radiation theory. To evaluate radiation fields generated by external sources, we must determine the particle velocity $\vec{v}$ and position $\vec{x}$ as functions of the lab-frame time $t$. As discussed earlier, Maxwell’s equations should be solved in the lab frame, incorporating the current and charge density produced by particles moving along their covariant trajectories. For arbitrary values of $v/c$, covariant calculations of radiation processes become highly complex. However, in certain cases, significant simplifications are possible. One such case is a non-relativistic radiation setup. The non-relativistic assumption enables the dipole approximation, which is of great practical importance. When considering only the dipole component of radiation, we disregard all details of the electron trajectory. Consequently, dipole radiation is insensitive to differences between covariant and non-covariant particle trajectories. However, this is merely the first and most practically relevant term in the expansion. To compute higher-order corrections beyond the dipole approximation, detailed knowledge of the electron trajectory is required. In such cases, relying on the non-covariant approach is insufficient, and the covariant trajectory must be used instead.
@@
It is generally assumed that Maxwell’s equations and the corrected form of Newton’s second law can explain all experiments conducted within a single inertial (laboratory) frame. According to standard textbooks, if only one frame is involved, knowledge of relativity is deemed unnecessary. Many physicists, having learned special relativity from textbooks, might say, “For those who want just enough understanding to solve problems, special relativity merely modifies Newton’s laws by introducing a correction factor to mass.” However, this is a misconception. The velocity of light transforms like that of a particle, and one cannot simply apply classical kinematics to mechanics while using Einstein’s kinematics for electrodynamics. We highlight errors in the standard approach to coupling fields and particles through a relatively simple example, ensuring that the core physical principles remain clear and unobscured by unnecessary mathematical complexity. We hope this example will draw greater attention to the crucial role of Wigner rotation in the transformation of non-collinear velocities.
@@
Accelerator physicists studying the application of relativity theory to synchrotron radiation often face a fundamental issue: the distinction between covariant and non-covariant particle trajectories has never been fully understood. Consequently, the contribution of relativistic kinematic effects to synchrotron radiation has been overlooked. Traditionally, accelerator physics has relied on Newtonian kinematics, which is incompatible with Maxwell’s equations. This raises an important question: if storage rings are designed without accounting for relativistic kinematics, how do they function? In reality, electron dynamics in storage rings are significantly influenced by radiation emission. We address this question in detail in Chapter 19, focusing on spontaneous synchrotron radiation from bending magnets and undulators.
@@
In Chapter 20, we critically reexamine existing XFEL theory, particularly for readers with limited knowledge of accelerator physics. We focus on the production of coherent undulator radiation by a modulated ultrarelativistic electron beam. Fortunately, understanding this process does not require prior knowledge of undulator radiation theory from Chapter 19, as it can be explained in a straightforward manner. Relativistic kinematics plays a fundamental role in XFEL physics, particularly through the rotation of the modulation wavefront, which, in the ultrarelativistic approximation, is closely linked to the relativity of simultaneity. When particle trajectories are analyzed in a Lorentz reference frame—an inertial frame where time coordinates are assigned using the Einstein synchronization procedure—relativistic kinematic effects, such as the relativity of simultaneity, must be considered. In the ultrarelativistic limit, the orientation of the modulation wavefront, i.e. the orientation of the plane of simultaneity, is always perpendicular to the electron beam velocity when the evolution of the beam’s evolution is described using Lorentz coordinates.
It is important to note that Maxwell’s equations are valid only in Lorentz reference frames. Einstein’s time order must be consistently applied in both dynamics and electrodynamics. Maxwell’s equations dictate that the wavefront of a laser beam is always orthogonal to its propagation direction. At this point, we emphasize that, according to relativity, a modulated electron beam in the ultrarelativistic asymptotic limit exhibits the same kinematics as a laser beam in Lorentz coordinates. Experiments show that this prediction is, in fact, true.
@@ 2.1 What is Special Relativity?
The laws of physics are invariant under Lorentz transformations. This invariance is a restrictive principle—it does not, by itself, determine the precise form of the dynamics involved. Understanding the postulates of special relativity is analogous to grasping the principle of energy conservation: initially, we accept it as a foundational principle, and later, we seek microscopic interpretations that are consistent with it. For any system governed by energy conservation, there should exist a more fundamental theory that provides insight into the detailed physical processes. Naturally, this deeper theory must also obey the principle of energy conservation.
The conservation of energy is a powerful tool, allowing us to analyze systems even in the absence of a complete fundamental theory. A similar methodological perspective applies to the postulates of special relativity. For instance, suppose we do not understand the microscopic reasons for muon decay, but we do know the decay law in the muon’s rest frame. This decay law, in such a case, serves as a phenomenological model. Its relativistic generalization to any Lorentz frame allows us to predict, for example, the average distance a muon travels before decaying.
Consider the case of muon decay. Suppose we lack a detailed understanding of why a muon disintegrates but have empirical knowledge of its decay law in its rest frame. This law serves as a phenomenological model. By applying relativistic principles, we can extend this law to an arbitrary Lorentz frame, enabling us to predict the muon’s average travel distance. Specifically, if we apply a Lorentz transformation to the decay law, we find that after a distance $\gamma v\tau_{0}$, the muon population in the lab frame is reduced to half of its original value. Here, $\gamma$ is the relativistic factor, and $\tau_{0}$ is the proper lifetime. This result can be interpreted as an increase in the effective lifetime of the muon from $\tau_{0}$ to $\gamma\tau_{0}$ in the lab frame due to time dilation.
However, special relativity is not a complete theory. Constructive (or microscopic) theories, such as electrodynamics or quantum field theory, offer a deeper understanding of physical processes than purely restrictive frameworks like special relativity. Relativistic kinematics serves primarily as a descriptive tool for interpreting how matter fields appear to behave from the perspective of different observers. In principle, any relativistic quantity can be derived directly from the underlying theories of matter, without explicitly invoking relativity.
For example, muons exhibit relativistic behavior in motion because the field forces governing their decay obey Lorentz-covariant quantum field equations. In the ”microscopic” view of relativistic phenomena, the Lorentz covariance of the fundamental laws of physics remains, much like energy conservation, an empirical and unexplained fact—but ultimately, all explanations must terminate at some foundational level.
@@ 2.3 The Myth About the Incorrectness of Galilean Transformations
The role of Galilean transformations within the framework of relativity requires careful reconsideration. Many physicists continue to view Galilean transformations—specifically, transformations from an orthogonal Lorentz basis to a non-orthogonal one—as outdated and incorrect when applied to spatial coordinates and time. A common argument against their validity is that they fail to preserve the form-invariance of Maxwell’s equations under a change of inertial frame. This reasoning is frequently found in well-known textbooks and monographs  [^1] It is widely believed that Galilean transformations, being pre-relativistic, are incompatible with the special theory of relativity. For instance, Bohm [1] states: “… the Galilean law of addition of velocities implies that the speed of light should vary with the speed of the observing equipment. Since this predicted variation is contrary to fact, the Galilean transformations evidently cannot be correct.” Similar assertions appear in more recent pedagogical literature. Drake and Purvis [2], for example, write: “One of the great insights of relativity theory was the realization that Galilean transformations are incorrect. The proper way to translate space-time measurements between inertial frames is through Lorentz transformations.”.
However, the assertion that Galilean transformations are inherently incorrect is a misconception. Special relativity is fundamentally a theory of four-dimensional space-time with a pseudo-Euclidean geometry. From this perspective, the principle of relativity naturally emerges as a geometric property of space-time, which can be described using arbitrary coordinate systems  [^2] The mathematical claim that transitioning to arbitrary coordinates leaves the geometry of space-time unchanged is considered erroneous in standard textbooks. As L. Landau and E. Lifshitz note [3]: ”This formula is called the Galilean transformation. It is easy to verify that this transformation, as was to be expected, does not satisfy the requirements of the theory of relativity; it does not leave the interval between events invariant.” This misconception is attributed to a failure to distinguish between the convention-dependent and convention-invariant aspects of the theory. In pseudo-Euclidean geometry, the interval between events remains invariant under arbitrary coordinate transformations.. Thus, contrary to the standard presentation in most textbooks, Galilean transformations are in fact compatible with the principle of relativity, even though they modify the explicit form of Maxwell’s equations.
@@ 2.4 The Non-Relativistic Limit of Lorentz Transformations
We wish to emphasize the following point: An infinitesimal Lorentz transformation differs from the Galilean transformation only by incorporating the relativity of simultaneity. This is the sole relativistic effect that appears in the first order of $v/c$. Other higher-order effects, such as Lorentz-Fitzgerald contraction and time dilation, arise from higher-order terms and can be derived mathematically by iterating the infinitesimal transformation  [^3] There is a common misconception that the Lorentz transformation reduces to the Galilean transformation in the non-relativistic limit. For example, French [4] states, ”The reduction of $t^{\prime}=\gamma(t-vx/c^{2})$ to the Galilean relation $t^{\prime}=t$ requires $x\ll ct$ as well as $v/c\ll 1$”. Similar statements are found in recent textbooks, such as Rafelski [5], who asserts: ”A wealth of daily experience shows that the Galilean coordinate transformation (GT) is correct in the nonrelativistic limit in which the speed of light is so large that it plays no physical role. Any coordinate transformation replacing the GT must also agree with this experience, and thus must contain the GT in the nonrelativistic limit.” This conclusion is mathematically flawed. As Baierlein [6] points out: ”If the Lorentz transformation for infinitesimal $v/c$ were to reduce to the Galilean transformation, then the iterative process could never generate a finite Lorentz transformation that is radically different from the Galilean transformation..
@@ 2.6 Convention-Dependent Aspects of the Theory
Now, if we take the origin of a Cartesian coordinate system and connect it to a point on the path, the resulting vector is the position vector $\vec{x}(l)$  [^6] For a more detailed discussion on the distinction between path and trajectory, see [10].. The key difference between trajectory $\vec{x}(t)$ and path $\vec{x}(l)$ is fundamental: the path has an exact, objective meaning—it is independent of any convention. In contrast, and consistent with the inherent conventionality of velocity, the trajectory $\vec{x}(t)$ is convention-dependent and lacks an exact objective interpretation.
@@ 2.8 Relativistic Particle Dynamics
This solution to the dynamics problem in the lab frame does not rely on Lorentz transformations. Conventional particle tracking treats space-time in a non-relativistic manner, as a (3+1) manifold. In this approach, the only modification to classical mechanics is the introduction of relativistic mass, while time remains distinct from space. There is no mixing of spatial and temporal coordinates  [^7] Clarification of the true content of the non-covariant theory can be found in various advanced textbooks. To quote e.g. Ferrarese and Bini [12]: ”… within a single inertial frame, the time is an absolute quantity in special relativity also. As a consequence, if no more than one frame is involved, one would not expect differences between classical and relativistic kinematics. But in the relativistic context, there are differences in the transformation laws of the various relative quantities (of kinematics or dynamics), when passing from one reference frame to another.” These authors emphasize the unique role of a ”single inertial frame.” This term suggests that a defining characteristic of non-covariant theory is the absence of relativistic kinematics in describing particle motion. A similar viewpoint is expressed by Friedman [15]: ”Within any single inertial frame, things look precisely the same as in Newtonian kinematics: there is an enduring Euclidean three-space, a global time $t$, and the law of motion.”.

### v12

@@
Current draft was completed by a new Chapter 17. If you have any comments, please send an email to evgueni.saldin@desy.de.
@@ 1 Introduction
Solving problems related to covariant particle tracking presents two key challenges. First, conventional particle tracking within a single inertial frame is typically based on classical Newtonian mechanics, which does not incorporate Lorentz transformations. Second, there is considerable confusion regarding the combination of non-parallel velocities in special relativity. The standard textbook treatment of relativistic velocity transformation implicitly assumes that the $(x^{\prime},y^{\prime},z^{\prime})$ axes of a moving observer remain parallel to the $(x,y,z)$ axes of the lab frame. In other words, it presumes that both observers share a common three-dimensional space—a misconception. Textbooks often overlook the fact that in Lorentz coordinatization, two observers following different trajectories experience different three-spaces. Special relativity lacks an absolute notion of simultaneity, leading to a fundamental mixing of spatial and temporal coordinates. For one observer, spatial measurements inherently include a contribution from the time coordinate as perceived by another observer. Consequently, there is no well-defined concept of an instantaneous three-space.
@@
Our analysis highlights the distinction between the concepts of path and trajectory. A path is a purely geometric notion, representing the spatial course of a particle without reference to time. In contrast, a trajectory provides more information, as it specifies not only the particle’s spatial position but also the corresponding time instant.
@@
The path has a precise, objective meaning. For example, in a magnetic field, the curvature radius of the path—and consequently, the three-momentum—is convention-invariant and thus objectively defined. However, the trajectory $\vec{x}(t)$ is convention-dependent, reflecting the inherent conventionality of velocity, and therefore lacks an exact objective meaning. While dynamical theory involves the concept of a particle’s trajectory, direct verification of it is unnecessary. Instead, the trajectory serves as a theoretical tool in the analysis of electrodynamics problems.
In Chapter 16, we derived specific quantitative results for relativistic velocity addition using only elementary principles of special relativity. However, readers seeking a more mathematically rigorous treatment may prefer Chapter 17, where we present a comprehensive account of the mathematical structure underlying the composition of Lorentz boosts.
@@
In Chapter 19, we present a critical reexamination of radiation theory. To evaluate radiation fields generated by external sources, we must determine the particle velocity $\vec{v}$ and position $\vec{x}$ as functions of the lab-frame time $t$. As discussed earlier, Maxwell’s equations should be solved in the lab frame, incorporating the current and charge density produced by particles moving along their covariant trajectories. For arbitrary values of $v/c$, covariant calculations of radiation processes become highly complex. However, in certain cases, significant simplifications are possible. One such case is a non-relativistic radiation setup. The non-relativistic assumption enables the dipole approximation, which is of great practical importance. When considering only the dipole component of radiation, we disregard all details of the electron trajectory. Consequently, dipole radiation is insensitive to differences between covariant and non-covariant particle trajectories. However, this is merely the first and most practically relevant term in the expansion. To compute higher-order corrections beyond the dipole approximation, detailed knowledge of the electron trajectory is required. In such cases, relying on the non-covariant approach is insufficient, and the covariant trajectory must be used instead.
@@
It is generally assumed that Maxwell’s equations and the corrected form of Newton’s second law can explain all experiments conducted within a single inertial (laboratory) frame. According to standard textbooks, if only one frame is involved, knowledge of relativity is deemed unnecessary. Many physicists, having learned special relativity from textbooks, might say, “For those who want just enough understanding to solve problems, special relativity merely modifies Newton’s laws by introducing a correction factor to mass.” However, this is a misconception. The velocity of light transforms like that of a particle, and one cannot simply apply classical kinematics to mechanics while using relativistic kinematics for electrodynamics. We highlight errors in the standard approach to coupling fields and particles through a relatively simple example, ensuring that the core physical principles remain clear and unobscured by unnecessary mathematical complexity. We hope this example will draw greater attention to the crucial role of Wigner rotation in the transformation of non-collinear velocities.
@@
Accelerator physicists studying the application of relativity theory to synchrotron radiation often face a fundamental issue: the distinction between covariant and non-covariant particle trajectories has never been fully understood. Consequently, the contribution of relativistic kinematic effects to synchrotron radiation has been overlooked. Traditionally, accelerator physics has relied on Newtonian kinematics, which is incompatible with Maxwell’s equations. This raises an important question: if storage rings are designed without accounting for relativistic kinematics, how do they function? In reality, electron dynamics in storage rings are significantly influenced by radiation emission. We address this question in detail in Chapter 20, focusing on spontaneous synchrotron radiation from bending magnets and undulators.
@@
In Chapter 21, we critically reexamine existing XFEL theory, particularly for readers with limited knowledge of accelerator physics. We focus on the production of coherent undulator radiation by a modulated ultrarelativistic electron beam. Fortunately, understanding this process does not require prior knowledge of undulator radiation theory from Chapter 20, as it can be explained in a straightforward manner. Relativistic kinematics plays a fundamental role in XFEL physics, particularly through the rotation of the modulation wavefront, which, in the ultrarelativistic approximation, is closely linked to the relativity of simultaneity. When particle trajectories are analyzed in a Lorentz reference frame—an inertial frame where time coordinates are assigned using the Einstein synchronization procedure—relativistic kinematic effects, such as the relativity of simultaneity, must be considered. In the ultrarelativistic limit, the orientation of the modulation wavefront, i.e. the orientation of the plane of simultaneity, is always perpendicular to the electron beam velocity when the evolution of the beam’s evolution is described using Lorentz coordinates.
It is important to recognize that Maxwell’s equations are valid only within Lorentz reference frames. Consequently, Einstein’s concept of time ordering must be applied consistently in both dynamics and electrodynamics. According to Maxwell’s equations, the wavefront of a laser beam is always orthogonal to its direction of propagation. Notably, in the framework of special relativity, a modulated electron beam in the ultrarelativistic (asymptotic) limit exhibits the same kinematic behavior as a laser beam when described in Lorentz coordinates. Experimental evidence confirms the accuracy of this prediction.
@@ 2.1 What is Special Relativity?
Special relativity is founded on the principle that the laws of physics are invariant under Lorentz transformations. This principle is inherently restrictive: it does not dictate the specific form of the dynamics involved, but rather constrains any physical theory to be consistent with Lorentz symmetry.
Understanding the postulates of special relativity is conceptually similar to understanding the principle of energy conservation. Initially, we encounter energy conservation as a general rule. Later, we seek deeper, microscopic theories—such as classical mechanics or quantum field theory—that not only comply with this principle but also explain the underlying mechanisms. These deeper theories must, of course, recover energy conservation as a consequence of their more fundamental laws.
The utility of the energy conservation principle lies in its ability to guide analysis, even when the full fundamental theory is not yet known. A similar methodological perspective applies to the postulates of special relativity. For example, suppose we do not fully understand the microscopic mechanism behind muon decay. However, if we know the decay law in the muon’s rest frame, we can treat this as a phenomenological law. By applying the principles of special relativity, we can generalize this decay law to any Lorentz frame and make testable predictions.
In particular, if we transform the decay law to the laboratory frame using a Lorentz boost, we find that the population of muons is reduced to half its original value after they travel a distance of $\gamma v\tau_{0}$, where $\gamma$ is the Lorentz factor and $\tau_{0}$ is the muon’s proper lifetime. This implies that, from the lab’s perspective, the muon’s lifetime appears to be extended to $\gamma\tau_{0}$. Such predictions are direct consequences of the relativistic transformation of time intervals.
Despite its successes, special relativity is not a complete theory. It is a restrictive framework—more a set of symmetry constraints than a constructive (microscopic) explanation. Constructive theories, such as electrodynamics or quantum field theory, provide a more detailed understanding of physical phenomena. Special relativity, in this context, serves primarily as a kinematical interpretation of how physical processes appear from different inertial frames.
Importantly, all relativistic effects can, in principle, be derived directly from underlying physical laws—without invoking relativity as an independent framework—provided those laws are Lorentz covariant. For instance, muons in motion exhibit relativistic behavior because the quantum field equations governing their interactions and decay are Lorentz invariant.
In the ”microscopic” view of relativistic phenomena, the Lorentz covariance of the fundamental laws of physics remains, much like energy conservation, an empirical and unexplained fact—but ultimately, all explanations must terminate at some foundational level.
@@ 2.3 The Myth About the Incorrectness of Galilean Transformations
The role of Galilean transformations within the framework of relativity requires careful reconsideration. Many physicists continue to view Galilean transformations—specifically, transformations from an orthogonal Lorentz basis to a non-orthogonal one—as outdated and incorrect when applied to spatial coordinates and time. A common argument against their validity is that they fail to preserve the form-invariance of Maxwell’s equations under a change of inertial frame. This reasoning is frequently found in well-known textbooks and monographs.  [^1] It is widely believed that Galilean transformations, being pre-relativistic, are incompatible with the special theory of relativity. For instance, Bohm [1] states: “… the Galilean law of addition of velocities implies that the speed of light should vary with the speed of the observing equipment. Since this predicted variation is contrary to fact, the Galilean transformations evidently cannot be correct.” Similar assertions appear in more recent pedagogical literature. Drake and Purvis [2], for example, write: “One of the great insights of relativity theory was the realization that Galilean transformations are incorrect. The proper way to translate space-time measurements between inertial frames is through Lorentz transformations.”
However, the assertion that Galilean transformations are inherently incorrect is a misconception. Special relativity is fundamentally a theory of four-dimensional space-time with a pseudo-Euclidean geometry. From this perspective, the principle of relativity naturally emerges as a geometric property of space-time, which can be described using arbitrary coordinate systems.  [^2] The mathematical claim that transitioning to arbitrary coordinates leaves the geometry of space-time unchanged is considered erroneous in standard textbooks. As L. Landau and E. Lifshitz note [3]: ”This formula is called the Galilean transformation. It is easy to verify that this transformation, as was to be expected, does not satisfy the requirements of the theory of relativity; it does not leave the interval between events invariant.” This misconception is attributed to a failure to distinguish between the convention-dependent and convention-invariant aspects of the theory. In pseudo-Euclidean geometry, the interval between events remains invariant under arbitrary coordinate transformations. Thus, contrary to the standard presentation in most textbooks, Galilean transformations are in fact compatible with the principle of relativity, even though they modify the explicit form of Maxwell’s equations.
@@ 2.4 The Non-Relativistic Limit of Lorentz Transformations
We wish to emphasize the following point: An infinitesimal Lorentz transformation differs from the Galilean transformation only by incorporating the relativity of simultaneity. This is the sole relativistic effect that appears in the first order of $v/c$. Other higher-order effects, such as Lorentz-Fitzgerald contraction and time dilation, arise from higher-order terms and can be derived mathematically by iterating the infinitesimal transformation.  [^3] There is a common misconception that the Lorentz transformation reduces to the Galilean transformation in the non-relativistic limit. For example, French [4] states, ”The reduction of $t^{\prime}=\gamma(t-vx/c^{2})$ to the Galilean relation $t^{\prime}=t$ requires $x\ll ct$ as well as $v/c\ll 1$”. Similar statements are found in recent textbooks, such as Rafelski [5], who asserts: ”A wealth of daily experience shows that the Galilean coordinate transformation (GT) is correct in the nonrelativistic limit in which the speed of light is so large that it plays no physical role. Any coordinate transformation replacing the GT must also agree with this experience, and thus must contain the GT in the nonrelativistic limit.” This conclusion is mathematically flawed. As Baierlein [6] points out: ”If the Lorentz transformation for infinitesimal $v/c$ were to reduce to the Galilean transformation, then the iterative process could never generate a finite Lorentz transformation that is radically different from the Galilean transformation.”
@@ 2.6 Convention-Dependent Aspects of the Theory
Now, if we take the origin of a Cartesian coordinate system and connect it to a point on the path, the resulting vector is the position vector $\vec{x}(l)$.  [^6] For a more detailed discussion on the distinction between path and trajectory, see [10]. The key difference between trajectory $\vec{x}(t)$ and path $\vec{x}(l)$ is fundamental: the path has an exact, objective meaning—it is independent of any convention. In contrast, and consistent with the inherent conventionality of velocity, the trajectory $\vec{x}(t)$ is convention-dependent and lacks an exact objective interpretation.
@@ 2.8 Relativistic Particle Dynamics
This solution to the dynamics problem in the lab frame does not rely on Lorentz transformations. Conventional particle tracking treats space-time in a non-relativistic manner, as a (3+1) manifold. In this approach, the only modification to classical mechanics is the introduction of relativistic mass, while time remains distinct from space. There is no mixing of spatial and temporal coordinates.  [^7] Clarification of the true content of the non-covariant theory can be found in various advanced textbooks. To quote e.g. Ferrarese and Bini [12]: ”… within a single inertial frame, the time is an absolute quantity in special relativity also. As a consequence, if no more than one frame is involved, one would not expect differences between classical and relativistic kinematics. But in the relativistic context, there are differences in the transformation laws of the various relative quantities (of kinematics or dynamics), when passing from one reference frame to another.” These authors emphasize the unique role of a ”single inertial frame.” This term suggests that a defining characteristic of non-covariant theory is the absence of relativistic kinematics in describing particle motion. A similar viewpoint is expressed by Friedman [15]: ”Within any single inertial frame, things look precisely the same as in Newtonian kinematics: there is an enduring Euclidean three-space, a global time $t$, and the law of motion.”